# MultiAPI Spoof: A Multi-API Dataset and Local-Attention Network for Speech Anti-spoofing Detection thanks: Corresponding Author: Ming Li

### v2

@@ Abstract
Existing speech anti-spoofing benchmarks rely on a narrow set of public models, creating a substantial gap from real-world scenarios in which commercial systems employ diverse, often proprietary APIs. To address this issue, we introduce MultiAPI Spoof, a multi-API audio anti-spoofing dataset comprising about 230 hours of synthetic speech generated by 30 distinct APIs, including commercial services, open-source models, and online platforms. Based on this dataset, we define the API tracing task, enabling fine-grained attribution of spoofed audio to its generation source. We further propose Nes2Net-LA, a local-attention enhanced variant of Nes2Net that improves local context modeling and fine-grained spoofing feature extraction. Experiments show that Nes2Net-LA achieves state-of-the-art performance and offers superior robustness, particularly under diverse and unseen spoofing conditions. Code  [^1] https://github.com/XuepingZhang/MultiAPI-Spoof and dataset  [^2] https://xuepingzhang.github.io/MultiAPI-Spoof-Dataset/ have released.
@@ 1 Introduction
Recently, Text-To-Speech (TTS) [tts1, tts2, tts3, tts4], Voice Conversion (VC) [vc1, vc2, vc3, vc4], and generative modeling techniques [qwen2audio, gpa, gpt4, Gemini] have evolved rapidly. In particular, end-to-end dialogue systems [fun, Qwen3-omni, wavbench, Baichuan-omni], speech continuation models [Baichuan-audio], and style- or emotion-specific speech generation models [SoulXPodcast, flexivoice, Step-Audio-EditX] have advanced significantly. As a result, synthetic speech has become increasingly realistic and pervasive in everyday applications. Modern audio generation systems, especially those based on diffusion and large-scale generative models [diffu, Qwen3-omni, Kimi-audio], can now produce speech that closely mimics human prosody, timbre, and emotion. However, they have introduced serious security risks and can be easily misused for impersonation or misinformation.
@@
    Furthermore, we propose a new detection method, namely Nes2Net-LA, built upon Nes2Net [nes2net]. By integrating local attention modules between Nested blocks, Nes2Net-LA enhances local context modeling and fine-grained spoofing feature extraction, thereby improving robustness and discriminative capability. The Nes2Net-LA achieves state-of-the-art (SOTA) performance across multiple anti-spoofing benchmarks.
@@ 2.2 Dataset Split
The MultiAPI Spoof dataset is partitioned by the APIs. APIs A0–A20 are used to construct the training, development, and evaluation subsets with a 70/10/20 split, ensuring sufficient variation within seen sources. APIs A21–A23 are reserved entirely for development, while APIs A24–A29 are held out exclusively for evaluation. This design enables two evaluation conditions:
@@ 3 Local Attention Enhanced Anti-spoofing Network
Figure 1: Overall architecture of the Nes2Net-X and proposed Nes2Net-LA frameworks. The model first extracts high-dimensional representations from the input audio and then processes them using nested multi-scale feature fusion. Nes2Net-LA further enhances cross-block interactions through a sliding-window local attention mechanism.
@@ 3.1 Prior Knowledge: Nested Res2Net (Nes2Net-X)
The Nes2Net-X architecture [nes2net], is a multi-scale feature extractor for high-dimensional speech representations. An audio segment $x_{i}$ is encoded into $x^{\prime}i\in\mathbb{R}^{C\times T^{\prime}}$ and split into channel-wise subsets $x_{i,1},\dots,x_{i,J}$. Each subset is processed hierarchically: the first passes through Convolution(Conv) and Weighted Summation (WS), while subsequent subsets are fused with the previous output before Conv. All outputs are refined via Conv and Squeeze-and-Excitation (SE) module with residual connections to obtain the anti-spoofing feature representations $h_{i,j}\in R^{(C/J)\times T^{\prime}}$, as shown in (1).
@@
wo MultiAPI Spoof XLSR+AASIST [xlsr+aasist] 2.02 / 0.026 / 0.029 - - 7.30 / 0.098 / 0.106 12.96 / 0.132 / 0.190
@@ 3.2 Nes2Net with Local Attention (Nes2Net-LA)
As shown in Figure 1, for each block, we define a local sliding-window neighborhood $\mathcal{N}(i,j)=\{h_{i,k}\mid k\in[j-K,j+K]\}$, where $K$ is the window radius. A local attention [att] operator is then applied to get the local feature representation $y_{i,j}\in\mathbb{R}^{(C/J)\times T^{\prime}}$, as shown in (3).
@@ 4 Anti-Spoofing API Tracing Task
Our baseline model uses hidden representations from the XLSR-300M [xlsr] encoder, followed by an attention pooling layer to aggregate embeddings from each encoder hidden layer, and ends with a squeeze-and-excitation (SE) layer to get the final results. During training, only the 21 seen APIs are used. At inference, samples whose maximum predicted probability falls below a threshold are classified as the unseen class, effectively turning the task into a 22-class classification problem. More details can be found in our open-source code [^3] https://github.com/XuepingZhang/MultiAPI-Spoof.
@@ 5.1 Experimental Setup
Dataset The anti-spoofing experiments are conducted on a collection of six public datasets: TIMIT [TIMIT], ODSS [ODSS], FoR [FoR], AI4T [AI4T], ASV5 [asv5], and MLAAD [MLAAD]. These corpora cover a wide variety of spoofing sources, including real-world collected data, text-to-speech (TTS), and voice conversion (VC). We consider two training configurations. In the first setting, the six datasets are merged into a single training set, and evaluation is performed across three target domains: ITW, MultiAPI Spoof test set, and AI4T test set. In the second setting, the MultiAPI Spoof training set is additionally included in the training pool, and the models are re-evaluated on the same three domains.
@@ 5.2.1 Anti-Spoofing on MultiAPI Spoof
In the first comparison experiment, models are trained on six commonly used anti-spoofing datasets (TIMIT [TIMIT], ODSS [ODSS], FoR [FoR], AI4T [AI4T], ASV5 [asv5], MLAAD [MLAAD]) without including MultiAPI Spoof training set. Two systems, XLSR+AASIST [xlsr+aasist] and XLSR+Nes2Net-X [nes2net], are evaluated on ITW [ITW], MultiAPI Spoof test set, and AI4T [AI4T], respectively. Both models present relatively high EERs on the MultiAPI Spoof evaluation set, indicating a domain shift that existing datasets fail to cover. In the second comparison experiment, MultiAPI Spoof training set is incorporated into the training pool. Across all evaluation sets, especially on the MultiAPI Spoof test set itself, both XLSR+AASIST and XLSR+Nes2Net-X show substantial reductions in EER, minDCF, and actDCF. For instance, XLSR+AASIST decreases the EER from 12.96% to 6.26% on MultiAPI Spoof test set, and similarly, XLSR+Nes2Net decreases the EER from 7.77% to 5.64%.
@@
Table 2: Comparison of our proposed XLSR+Nes2Net-LA system with recent state-of-the-art anti-spoofing models. “SP” denotes Sample Pruning [AI4T]; “RB” denotes RawBoost augmentation [Rawboost]; “C” denotes codec augmentation. “Data Collection 1” consists of ASVspoof 2019 [asv19], FoR[FoR], ASVspoof 2021 DF[asv21], TIMIT[TIMIT], ODSS[ODSS], MLAAD[MLAAD], and ASV5[asv5] training sets. “Data Collection 2” includes TIMIT, ODSS, FoR, AI4T [AI4T], ASV5, MLAAD, and the proposed MultiAPI Spoof training set.
@@ 5.2.2 Effectiveness of Local Attention (Nes2Net-LA)
As shown in Table 1 and Table 2, Nes2Net-LA trained on the data collection pool outperforms recent state-of-the-art models across all evaluation benchmarks, even without any data augmentation or pruning. The most substantial improvements are observed on the unseen split of the MultiAPI Spoof test set, demonstrating that the proposed local attention mechanism effectively enhances cross-block interactions and produces more discriminative and robust anti-spoofing representations.
@@ 5.2.3 API Tracing on MultiAPI Spoof
Table 3 presents the API tracing results on MultiAPI Spoof. The model is trained on the MultiAPI Spoof training set and tested on the dev and eval sets. We report results separately for seen and unseen API types. Overall performance is high across seen APIs, with both dev and eval achieving substantial precision, recall, and F1 scores. However, high precision but low recall for the unseen class shows that predictions are accurate, but many unseen-class instances are not correctly identified. This indicates that current methods for this task, particularly in handling unseen APIs, still require further investigation.
@@ 6 Conclusion
In this paper, we present MultiAPI Spoof, a multi-API speech anti-spoofing dataset, and first introduce the API tracing task for fine-grained source attribution. Experiments show that incorporating MultiAPI Spoof into training significantly improves cross-domain robustness. We also propose a local-attention enhanced anti-spoofing network, namely Nes2Net-LA. It consistently outperforms Nes2Net-X, achieving state-of-the-art performance, demonstrating its effectiveness in improving robustness and discriminative capability.

### v3

@@ Abstract
Existing speech anti-spoofing benchmarks rely on a narrow set of public models, creating a substantial gap from real-world scenarios in which commercial systems employ diverse, often proprietary APIs. To address this issue, we introduce MultiAPI Spoof, a multi-API audio anti-spoofing dataset comprising about 230 hours of synthetic speech generated by 30 distinct APIs, including commercial services, open-source models, and online platforms. Furthermore, we propose Nes2Net-LA, a local-attention enhanced variant of Nes2Net that improves local context modeling and fine-grained spoofing feature extraction. Based on this dataset, we also define the API tracing task, enabling fine-grained attribution of spoofed audio to its generation source. Experiments show that Nes2Net-LA achieves state-of-the-art performance and offers superior robustness, particularly under diverse and unseen spoofing conditions. Code  [^1] https://github.com/XuepingZhang/MultiAPI-Spoof and dataset  [^2] https://xuepingzhang.github.io/MultiAPI-Spoof-Dataset/ have been released.
@@ 1 Introduction
Recently, Text-To-Speech (TTS) [tts1, tts2, tts3, tts4], Voice Conversion (VC) [vc1, vc2, vc3, vc4], and generative modeling techniques [qwen2audio, gpa, gpt4, Gemini] have evolved rapidly. In particular, end-to-end dialogue systems [fun, Qwen3-omni, wavbench, Baichuan-omni], speech continuation models [Baichuan-audio, mimo-audio], and style- or emotion-specific speech generation models [SoulXPodcast, flexivoice, Step-Audio-EditX] have advanced significantly. As a result, synthetic speech has become increasingly realistic and pervasive in everyday applications. Modern audio generation systems, especially those based on diffusion and large-scale generative models [diffu, Qwen3-omni, Kimi-audio], can now produce speech that closely mimics human prosody, timbre, and emotion. However, they have introduced serious security risks and can be easily misused for impersonation or misinformation.
@@
    Furthermore, we propose a new anti-spoofing detection method, namely Nes2Net-LA, built upon Nes2Net [nes2net]. By integrating local attention modules between Nested blocks, Nes2Net-LA enhances local context modeling and fine-grained spoofing feature extraction, thereby improving robustness and discriminative capability. The Nes2Net-LA achieves state-of-the-art (SOTA) performance across multiple anti-spoofing benchmarks.
@@ 2.2 Dataset Split
The MultiAPI Spoof dataset is partitioned by the APIs. APIs A0–A20 are used to construct the training, development, and evaluation subsets with a 70/10/20 % split, ensuring sufficient variation within seen sources. APIs A21–A23 are reserved entirely for development, while APIs A24–A29 are held out exclusively for evaluation. This design enables two evaluation conditions:
@@ 3 Local Attention Enhanced Anti-spoofing Network
Figure 1: Overall architecture of proposed Nes2Net-LA frameworks. The model first extracts high-dimensional representations from the input audio and then processes them using nested multi-scale feature fusion. Nes2Net-LA further enhances cross-block interactions through a sliding-window local attention mechanism. `WS' represent Weighted Summation, and `ATT' represent scaled dot-product self-attention.
@@ 3.1 Prior Knowledge: Nested Res2Net (Nes2Net-X)
The Nes2Net-X architecture [nes2net] is a multi-scale feature extractor for high-dimensional speech representations. An audio segment $x_{i}$ is encoded into $x_{i}^{\prime}\in\mathbb{R}^{C\times T^{\prime}}$ and split into channel-wise subsets $x_{i,1},\dots,x_{i,J}$. Each subset is processed hierarchically: the first passes through Convolution(Conv) and Weighted Summation (WS), while subsequent subsets are fused with the previous output before convolution. All outputs are refined with a convolution and Squeeze-and-Excitation (SE) module with residual connections to obtain the anti-spoofing feature representations $h_{i,j}\in R^{(C/J)\times T^{\prime}}$, as shown in (1).
@@
without MultiAPI Spoof XLSR+AASIST [xlsr+aasist] 2.02 / 0.026 / 0.029 - - 7.30 / 0.098 / 0.106 12.96 / 0.132 / 0.190
@@ 3.2 Nes2Net with Local Attention (Nes2Net-LA)
As shown in Figure 1, for each block, we define a local sliding-window neighborhood $\mathcal{N}(i,j)=\{h_{i,k}\mid k\in[j-K,j+K]\}$, where $K$ is the window radius. A local scaled dot-product self-attention (ATT) [att] operator is then applied to get the local feature representation $y_{i,j}\in\mathbb{R}^{(C/J)\times T^{\prime}}$, as shown in (3).
@@ 4 Anti-Spoofing API Tracing Task
Our baseline model uses hidden representations from the XLSR-300M [xlsr] encoder, followed by an attention pooling layer to aggregate embeddings from each encoder hidden layer, and ends with a Squeeze-and-Excitation (SE) layer to get the final results. During training, only the 21 seen APIs are used. At inference, samples whose maximum predicted probability falls below a threshold are classified as the unseen class, effectively turning the task into a 22-class classification problem.
@@ 5.1 Experimental Setup
Dataset The anti-spoofing experiments are conducted on a collection of six public datasets: TIMIT [TIMIT], ODSS [ODSS], FoR [FoR], AI4T [AI4T], ASV5 [asv5], and MLAAD [MLAAD]. These corpora cover a wide variety of spoofing sources, including real-world collected data, text-to-speech (TTS), and voice conversion (VC). We consider two training configurations. In the first setting, the six datasets are merged into a single training set, and evaluation is performed across three target domains: the full ITW dataset [ITW], MultiAPI Spoof test set, and AI4T test set. In the second setting, the MultiAPI Spoof training set is additionally included in the training pool, and the models are re-evaluated on the same three domains.
@@ 5.2.1 Anti-Spoofing on MultiAPI Spoof
In the first comparison experiment, models are trained on six commonly used anti-spoofing datasets (TIMIT [TIMIT], ODSS [ODSS], FoR [FoR], AI4T [AI4T], ASV5 [asv5], MLAAD [MLAAD]) without including MultiAPI Spoof training set. Two systems, XLSR+AASIST [xlsr+aasist] and XLSR+Nes2Net-X [nes2net], are evaluated on ITW [ITW], MultiAPI Spoof test set, and AI4T [AI4T], respectively. Both models present relatively high EERs on the MultiAPI Spoof evaluation set, indicating a domain shift that existing datasets fail to cover. In the second comparison experiment, MultiAPI Spoof training set is incorporated into the training pool. Across all evaluation sets, especially on the MultiAPI Spoof test set itself, both XLSR+AASIST and XLSR+Nes2Net-X show substantial reductions in EER, minDCF, and actDCF. For instance, XLSR+AASIST decreases the EER from 7.30% to 0.70% on MultiAPI Spoof test set, and similarly, XLSR+Nes2Net decreases the EER from 7.08% to 0.69%.
@@
Table 2: Comparison of our proposed XLSR+Nes2Net-LA system with recent state-of-the-art anti-spoofing models. “SP” denotes Sample Pruning [AI4T]; “RB” denotes RawBoost augmentation [Rawboost]; “C” denotes codec augmentation. “Data Collection 1” consists of ASVspoof 2019 [asv19], FoR[FoR], ASVspoof 2021 DF[asv21], TIMIT[TIMIT], ODSS[ODSS], MLAAD[MLAAD], and ASV5[asv5] training sets. On top of Data Collection 1, Data Collection 2 replaces ASVspoof 2019 and ASVspoof 2021 DF with high-quality AI4T [AI4T] and the proposed MultiAPI Spoof training set.
@@ 5.2.2 Effectiveness of Local Attention (Nes2Net-LA)
As shown in Table 1 and Table 2, Nes2Net-LA trained on the data collection pool outperforms recent state-of-the-art models across all evaluation benchmarks, even without any data augmentation or pruning. The most substantial improvements are observed on the unseen split of the MultiAPI Spoof test set. These results demonstrate that the proposed local attention mechanism produces more discriminative and robust anti-spoofing representations.
@@ 5.2.3 API Tracing on MultiAPI Spoof
Table 3 presents the API tracing results on MultiAPI Spoof. The model is trained on the MultiAPI Spoof training set and tested on the dev and eval sets. We report results separately for seen and unseen API types. Overall performance is high across seen APIs, with both dev and eval achieving substantial precision, recall, and F1 scores. However, high precision but low recall for the unseen class shows that predictions are accurate, but many unseen-class instances are not correctly identified and are falsely rejected as unseen cases. This indicates that current methods for this task, particularly in handling unseen APIs, still require further investigation.
@@ 6 Conclusion
In this paper, we present MultiAPI Spoof, a multi-API speech anti-spoofing dataset, and further introduce the API tracing task for fine-grained source attribution. Experiments show that incorporating MultiAPI Spoof into training significantly improves cross-domain robustness. We also propose a local-attention enhanced anti-spoofing network, namely Nes2Net-LA. It outperforms Nes2Net-X, achieving state-of-the-art performance, demonstrating its effectiveness in improving robustness and discriminative capability.
7 Generative AI Use Disclosure
Large LanguageModels (LLMs) were used solely for manuscript polishing (e.g., rephrasing and grammar checks) to improve clarity and readability. The LLMs were not used for ideation, methodology, experimental design, data analysis, or result interpretation. All scientific content was produced and verified by the authors.